A rectangular box has dimensions 1 1/2 feet x 2 feet x 3 feet. What is the length of the longest object that can be put in the box, if the object can be placed in any position? This comes from the THEA Practice Test. Thanks
Answer by nerdybill(5404)

A rectangular box has dimensions 1 1/2 feet x 2 feet x 3 feet. What is the length of the longest object that can be put in the box, if the object can be placed in any position?
.
The diagonal of the 2x3 is:
x^2 = 2^2 + 3^2
x^2 = 4 + 9
x^2 = 13
x = sqrt(13)
.
The diagonal of the 1.5xsqrt(13) is:
x^2 = 1.5^2 + sqrt(13)^2
x^2 = 2.25 + 13
x^2 = 15.25
x = 3.91 inches